# AeroVerse: UAV-Agent Benchmark Suite for Simulating, Pre-training, Finetuning, and Evaluating Aerospace Embodied World Models

### v1

@@ VII-B Qualitative Analysis
From Figure 11, although the 3D-LLM [9] encodes the 3D environment and perceives its surroundings, it demonstrates limited generalization due to insufficient training on outdoor 3D urban data. When confronted with a 3D urban scene, the output of 3D-LLM[9] resembles a description of an indoor environment, leading to significant hallucinations. The findings indicate that the performance of these 2D visual-language models surpasses that of 3D-LLM [9]. This superiority can be attributed to the greater number of training image-text pairs available for the 2D visual-language models, which enhances their generalization capabilities. Furthermore, they deliver more accurate descriptions based on egocentric view images of urban settings. However, instances of hallucinations persist, such as the erroneous description of a fire hydrant in front of a building with windows in instructblip-vicuna-7b [38], despite the absence of a fire hydrant in the image. Similarly, llama-adapter-v2-7B [39] inaccurately describes an individual walking in the distance, even though no person is present in the image.
@@
In the example illustrated in the left of Figure 12, the 3D visual-language model demonstrates the capability to perform short-term spatial reasoning based on its current posture and the 3D characteristics of the city to address questions. In contrast, the 2D visual-language model can derive answers solely from a single view image, which provides limited information. The results from this example indicate that the answer to the question is not present in the image. The building visible on the right reveals only a corner of the red roof in the bottom right of the image. This situation elucidates why 2D visual-language models, such as gpt-4o [20], instructblip-vicuna-7b [38], and qwen-lv-7b [42], frequently reference red buildings or state their inability to provide a definitive answer. Conversely, the input of the 3D-LLM [9] encompasses a more comprehensive and complex representation of urban features, leading to a correct and logical conclusion.
@@
In the example illustrated in the right of Figure 12, the model is tasked with following specific instructions to explore a distance before addressing subsequent questions. Due to the limitations of the 2D LLM, we derive its response to this question based on the parameters outlined in Section VI-A, effectively simplifying the complexity of the question. The answer provided by the 3D-LLM [9] is incorrect, as its input incorporates more complex 3D features. Currently, GPT-4-vision-preview and GPT-4o rank among the most advanced visual-language models, with GPT-4o demonstrating a slight advantage in addressing questions related to first-person view images. Instructblip-vicuna-7b [38] and llama-adapter-v2-7B [39] exhibit relatively weaker instruction-following capabilities, resulting in their inadequate responses to our inquiries. The responses from open-source 2D visual-language models, including blip2-flan-t5-xxl [35], llava-v1.6-vicuna-7b [32], and Mplug2 [41], are generally consistent with the gold standard, indicating their strong spatial reasoning and instruction-following abilities.
@@
As illustrated in Figure 13, each model exhibits distinct responses to this task. The 3D-LLM’s capability for indoor task planning does not generalize effectively to urban environments, resulting in answers that are inconsistent with our queries and resembling 3D scene captions instead. To address the limitations of the 2D visual-language model, we derive its response to this question based on the parameters outlined in Section VI-A. The Blip2-flan-t5-xxl model [35] fails to accurately represent the flight path as per our instructions, instead providing an interpretation similar to image captioning, which indicates a relatively poor ability to adhere to directives. In contrast, both gpt-4o and gpt-4-vision-review [20] deliver the most detailed and comprehensive analyses of the initial views along the trajectory. The Instructblip-vicuna-7b [38], llava-v1.6-vicuna-7b [32], Mplug2 [41], and llama-adapter-v2-7B [39] models do not describe the flight path in accordance with the timeline; instead, they provide a summary of the trajectory.
VII-C Discussion
@@
Scene Generalization Ability. To investigate the generalization ability of various models across different embodied scenes, we assessed the performance of all models on four tasks corresponding to their respective scenes and compared the average BLEU scores across these tasks, as illustrated in Figure 14 (a). In the campus scene, the dense buildings and numerous obstacles, such as trees, generally lead to poor performance from all models. Conversely, in the residential area, which is smaller and contains fewer objects, all models demonstrate improved performance. Among the models, qwen-lv-7b exhibits the best overall performance, achieving the highest scores in all four scenarios, closely followed by gpt-4o. In contrast, 3d-llm faces greater challenges due to the input being a three-dimensional scene rather than a modified image, resulting in subpar performance in each scenario.
@@
Task Generalization Ability. To investigate the generalization ability of various models across different embodied tasks, we evaluate the performance of all models in four scenarios based on the tasks, comparing the average BLEU scores as depicted in Figure 14 (b). The results indicate that InstructBLIP and BLIP2 excel in Task $1$, which exclusively assesses the models’ captioning capabilities. In contrast, the Llava, MiniGPT, and MPLUG series models demonstrate superior performance in Task $4$, which necessitates the integration, comprehension, and response to information.
@@
VIII Conclusion
To address the existing research gap in the aerospace embodied world model and to empower UAV intelligent agents with end-to-end autonomous perception, cognition, and action capabilities, we have developed a comprehensive benchmark suite named AeroVerse. This suite integrates simulation, pre-training, fine-tuning, and evaluation processes. Specifically, we establish a simulation platform, AeroSimulator, which encompasses four realistic urban scenarios tailored for the flight simulation of UAV intelligent agents. Furthermore, we introduce the first image-text pre-training dataset, AerialAgent-Ego10k, utilizing real drone footage captured from a first-person perspective, along with the virtual image-text-pose alignment dataset, CyberAgent-Ego500k, to facilitate the pre-training of aerospace embodied world models. To enhance the autonomous capabilities of UAV intelligent agents, we delineate five downstream tasks for the first time: 3D scene awareness, spatial reasoning, navigational exploration, task planning, and motion decision-making. In line with these tasks, we construct five instruction fine-tuning datasets: SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k, SkyAgent-Plan3k, and SkyAgent-Act3k. Additionally, we develop a downstream task evaluation method called SkyAgent-Eval, based on GPT-4, which reveals the potential and limitations of 2D/3D visual-language models in aerospace embodied tasks.
In the future, we plan to expand the simulation cities to include locations such as New York and to extend the scenarios to encompass forests, mountains, and other outdoor environments. We will continue to accumulate training datasets and refine the downstream tasks for UAV agents to complete the training of the aerospace embodied world model. This will subsequently promote applications in areas such as river garbage detection, power inspection, and forest fire rescue, thereby unlocking the application value of aerospace embodied intelligence.

### v2

@@ VIII-B Qualitative Analysis
From Figure 13, although the 3D-LLM [Hong20233DLLMIT] encodes the 3D environment and perceives its surroundings, it demonstrates limited generalization due to insufficient training on outdoor 3D urban data. When confronted with a 3D urban scene, the output of 3D-LLM[Hong20233DLLMIT] resembles a description of an indoor environment, leading to significant hallucinations. The findings indicate that the performance of these 2D visual-language models surpasses that of 3D-LLM [Hong20233DLLMIT]. This superiority can be attributed to the greater number of training image-text pairs available for the 2D visual-language models, which enhances their generalization capabilities. Furthermore, they deliver more accurate descriptions based on egocentric view images of urban settings. However, instances of hallucinations persist, such as the erroneous description of a fire hydrant in front of a building with windows in instructblip-vicuna-7b [Dai2023InstructBLIPTG], despite the absence of a fire hydrant in the image. Similarly, llama-adapter-v2-7B [Gao2023LLaMAAdapterVP] inaccurately describes an individual walking in the distance, even though no person is present in the image. By contrast, SkyAgent’s results align best with the ground truth: the road ahead branches to the left and right; on the right is a low-rise red rectangular building, behind in the distance is a low-rise pink rectangular building, and on the left is a low-rise blue rectangular building. At the same time, it does not mischaracterize the outdoor scene as an indoor one like 3D‑LLM does.
@@
In the example illustrated in the left of Figure 14, the 3D visual-language model demonstrates the capability to perform short-term spatial reasoning based on its current posture and the 3D characteristics of the city to address questions. In contrast, the 2D visual-language model can derive answers solely from a single view image, which provides limited information. The results from this example indicate that the answer to the question is not present in the image. The building visible on the right reveals only a corner of the red roof in the bottom right of the image. This situation elucidates why 2D visual-language models, such as GPT-4o [Achiam2023GPT4TR], instructblip-vicuna-7b [Dai2023InstructBLIPTG], and qwen-lv-7b [Bai2023QwenVLAV], frequently reference red buildings or state their inability to provide a definitive answer. Conversely, the input of the 3D-LLM [Hong20233DLLMIT] encompasses a more comprehensive and complex representation of urban features, leading to a correct and logical conclusion. In this case, both SkyAgent and SkyAgentX produce the correct answer, “white.” SkyAgent uses an embodied chain-of-thought to parse the constraints—“go straight, pass on the right, reach the airspace above the white building”—suppressing the distraction from the red roof and determining from relative orientation and path relations that the passed building is white. Building on this, SkyAgentX introduces multitask curriculum learning to align perception, reasoning, navigation, and planning, achieving more robust constraint propagation and a more concise explanation, clearly outperforming 2D models.
@@
In the example illustrated in the right of Figure 14, the model is tasked with following specific instructions to explore a distance before addressing subsequent questions. Due to the limitations of the 2D LLM, we derive its response to this question based on the parameters outlined in Section VI-A, effectively simplifying the complexity of the question. The answer provided by the 3D-LLM [Hong20233DLLMIT] is incorrect, as its input incorporates more complex 3D features. Currently, GPT-4-vision-preview and GPT-4o rank among the most advanced visual-language models, with GPT-4o demonstrating a slight advantage in addressing questions related to first-person view images. Instructblip-vicuna-7b [Dai2023InstructBLIPTG] and llama-adapter-v2-7B [Gao2023LLaMAAdapterVP] exhibit relatively weaker instruction-following capabilities, resulting in their inadequate responses to our inquiries. The responses from open-source 2D visual-language models, including blip2-flan-t5-xxl [Li2023BLIP2BL], llava-v1.6-vicuna-7b [Liu2023ImprovedBW], and Mplug2 [Ye2023mPLUGOwl2RM], are generally consistent with the gold standard, indicating their strong spatial reasoning and instruction-following abilities. In this case, SkyAgent outputs “One” which matches the ground truth. Its reasoning first anchors the blue-and-white irregular mid-rise building on the left, then infers via “turn left—go straight—turn right—count.” Benefiting from the aerospace embodied chain-of-thought and instruction fine-tuning, SkyAgent shows strong instruction-following ability, but references geometric details in first-person images insufficiently, which may cause ambiguity in complex alley scenes. With the gains from multitask curriculum learning, SkyAgentX is more robust under viewpoint changes and variations in target visibility, reducing hallucinations in intersection counting and spatial orientation, and overall performs better than mainstream 2D models.
@@
As illustrated in Figure 15, each model exhibits distinct responses to this task. The 3D-LLM’s capability for indoor task planning does not generalize effectively to urban environments, resulting in answers that are inconsistent with our queries and resembling 3D scene captions instead. To address the limitations of the 2D visual-language model, we derive its response to this question based on the parameters outlined in Section VI-A. The Blip2-flan-t5-xxl model [Li2023BLIP2BL] fails to accurately represent the flight path as per our instructions, instead providing an interpretation similar to image captioning, which indicates a relatively poor ability to adhere to directives. In contrast, both GPT-4o and GPT-4-vision-review [Achiam2023GPT4TR] deliver the most detailed and comprehensive analyses of the initial views along the trajectory. The Instructblip-vicuna-7b [Dai2023InstructBLIPTG], llava-v1.6-vicuna-7b [Liu2023ImprovedBW], Mplug2 [Ye2023mPLUGOwl2RM], and llama-adapter-v2-7B [Gao2023LLaMAAdapterVP] models do not describe the flight path in accordance with the timeline; instead, they provide a summary of the trajectory. SkyAgent can produce a step-wise flight path aligned with the timeline, anchoring the route using landmarks such as parking lot boundaries and black–white/blue–brown buildings, with markedly better instruction adherence than 2D/3D baselines. However, its semantic granularity is coarse, occasional left–right orientation drift occurs, and it does not explicitly bind the given poses and obstacle-avoidance constraints. With multitask curriculum learning, SkyAgentX further improves timeline alignment and landmark consistency; its step count and turns more closely match the actual trajectory, terminal target identification is more accurate, and hallucinations are reduced. Nonetheless, its expression of geometric constraints such as altitude and speed remains insufficient, and some environmental elements are described redundantly.
IX Discussion
@@
Scene Generalization Ability. To investigate the generalization ability of various models across different embodied scenes, we assessed the performance of all models on four tasks corresponding to their respective scenes and compared the average BLEU scores across these tasks, as illustrated in Figure 16 (a). In the campus scene, the dense buildings and numerous obstacles, such as trees, generally lead to poor performance from all models. Conversely, in the residential area, which is smaller and contains fewer objects, all models demonstrate improved performance. Among the models, our proposed SkyAgent and SkyAgentX exhibit the best overall performance, significantly outperforming all other models in all four scenarios. In contrast, 3d-llm faces greater challenges due to the input being a three-dimensional scene rather than a modified image, resulting in subpar performance in each scenario.
@@
Task Generalization Ability. To investigate the generalization ability of various models across different embodied tasks, we evaluate the performance of all models in four scenarios based on the tasks, comparing the average BLEU scores as depicted in Figure 16 (b). The results indicate that InstructBLIP and BLIP2 excel in Task $1$, which exclusively assesses the models’ captioning capabilities. In contrast, the Llava, MiniGPT, and MPLUG series models demonstrate superior performance in Task $4$, which necessitates the integration, comprehension, and response to information. Notably, our proposed SkyAgent and SkyAgentX models achieve state-of-the-art performance across all four tasks. They not only demonstrate exceptional captioning abilities in Task $1$ but also show superior capabilities in the more complex tasks (Tasks $2$, 33, and $4$) that require reasoning, navigation, and planning, surpassing the other model series.
@@
Technology Ethics. The potential risks of the autonomous drone technology involved in this study mainly include:
1. Privacy leakage risk: Visual sensors equipped on drones may inadvertently capture personal identification information (such as license plates or faces), requiring strict desensitization during the data collection phase.
2. Airspace safety issues: If autonomous flight algorithms are exploited maliciously, they could lead to collision accidents. We have built no-fly zone settings into AeroSimulator (drones must not fly below 30m or above 50m).
Scalability and System Requirements. To assess the scalability of our system across diverse and complex 3D urban environments, we conduct experiments under different hardware configurations. The platform supports the loading of large-scale realistic 3D city scenes, with its upper bound determined primarily by available GPU memory and computational power. For instance, the Shenzhen scene can be smoothly rendered on a GPU with 6GB VRAM (e.g., NVIDIA GTX 1060), while other more complex scenes require higher-end hardware such as an NVIDIA RTX 3090 for stable operation. These results demonstrate the system’s flexibility and adaptability to varying performance constraints.
Novelty and Comparative Analysis. This work’s primary novelty lies in establishing AeroVerse as the first comprehensive benchmark suite for aerospace embodied intelligence, addressing a critical gap in the literature. While powerful simulation tools like AirSim exist, they function as underlying engines rather than integrated benchmarks. The core research gap has been the absence of a full-stack platform that includes standardized task definitions, large-scale and domain-specific datasets, and unified evaluation protocols. AeroVerse is designed to fill this void. In comparison to recent related work such as OpenFly, AeroVerse offers several distinct contributions. Firstly, it is made public nearly six months prior, establishing its temporal precedence. Secondly, its scope is broader; whereas OpenFly concentrates on vision-language navigation, AeroVerse defines a five-dimensional task framework encompassing scene awareness, spatial reasoning, navigation, task planning, and motion decision. Thirdly, it introduces innovative data types, including the first large-scale real-world image-text pre-training dataset from a UAV perspective (AerialAgent-Ego15k), a virtual image-text-pose alignment dataset (CyberAgent-Ego500k), and five expert-labelled high-quality instruction datasets for downstream aerospace embodied tasks, which are crucial for pre-training and finetuning robust aerospace world models. Additionally, AeroVerse contains SkyAgentX, which pioneers as the first UAV-agent embodied large model, unifying “perception-reasoning-navigating-planning” into an end-to-end framework by integrating aerospace embodied chain-of-thought and multitask curriculum learning. Finally, our benchmark provides key scientific insights by systematically evaluating current large 2D/3D visual-language models, revealing their common failure modes in complex aerospace embodied tasks and highlighting the need for specialized aerospace embodied world models.
Bridging Simulation to the Real World. A core design principle of AeroVerse is to facilitate the transfer of models from simulation to real-world applications (i.e., Sim-to-Real). The benchmark bridges this gap by integrating a large-scale real-world dataset (AerialAgent-Ego15k) with a high-fidelity simulated dataset (CyberAgent-Ego500k). This dual-pronged approach provides models with a training foundation that combines real-world perceptual capabilities with the diverse interactive learning opportunities available in simulation. To validate this approach, we have conducted preliminary real-world deployments. An aerospace embodied world model trained on the AeroVerse benchmark is successfully deployed on a physical UAV, achieving autonomous perception and navigation in real urban scenarios. This result provides initial evidence that models trained within our benchmark possess the potential to transfer effectively to real-world applications and can serve as a strong foundation for future development in areas such as urban inspection and logistics delivery.
Limitations and Future Work. While this study makes significant contributions, we acknowledge several limitations that pave the way for future research.
1. Regarding the generalization of 3D models, our findings indicate that existing models like 3D-LLM exhibit weak generalization in complex urban scenes, largely because they are primarily trained on indoor data and struggle with analyzing complex outdoor spatial relationships. Future work should focus on two key directions: 1) Domain Knowledge Injection, which involves encoding prior knowledge from urban planning (e.g., building height distributions, road hierarchies) as constraints to guide model learning, and 2) Enhanced Spatial Relationship Pretraining through dedicated pretraining phases to deepen the model’s understanding of object relations in 3D space.
2. The diversity of simulation scenes is currently limited to four urban environments. Although representative, future iterations of AeroSimulator should incorporate more varied geographical settings, such as rural, mountainous, and coastal areas, to improve model robustness. We intend to integrate procedural generation techniques to dynamically create a wider range of scenes. We have already expanded our platform to include 20 high-fidelity 3D real-world urban scenes, significantly increasing the benchmark’s richness and the potential for model generalization.
3. The scale of real-world data, while pioneering, can be expanded. Although AerialAgent-Ego15k is the first large-scale dataset of its kind, we plan to continuously augment it with more data covering diverse weather conditions, lighting scenarios, and dynamic events to further close the Sim-to-Real gap.
X Conclusion
This paper addresses the research gap in aerospace embodied world models by developing AeroVerse, a comprehensive benchmark suite designed to empower UAV intelligent agents with end-to-end autonomous capabilities. We establish the AeroSimulator platform with four realistic urban scenarios, introduced two pioneering pre-training datasets (the real-world AerialAgent-Ego15k and the virtual CyberAgent-Ego500k), clearly delineated five standardized downstream tasks for the first time, and constructed corresponding instruction fine-tuning datasets. Our extensive experiments systematically reveal, for the first time, the widespread limitations of current mainstream 2D and 3D visual-language models in aerospace embodied tasks, highlighting critical areas for improvement such as spatial reasoning and long-term planning. Furthermore, we propose SkyAgentX, the first UAV-agent embodied large model integrating “perception-reasoning-navigating-planning”, which demonstrates superior performance, thus validating the effectiveness of our benchmark and providing a strong baseline for future models.
The significance of this study lies in providing the research community with a standardized platform for fair comparison and rapid iteration. The release of AeroVerse, along with our findings, will direct stakeholders and researchers by offering valuable references and a clear pathway for developing more robust, capable, and safer aerospace intelligent agents. In the future, we plan to expand the simulation environments and continue to refine the training datasets and downstream tasks, promoting applications in areas such as river garbage detection, power inspection, and forest fire rescue, thereby unlocking the full application value of aerospace embodied intelligence.
[Uncaptioned image] Fanglong Yao received the B.Sc. degree from Inner Mongolia University, Hohhot, China, in 2017, and the Ph.D. degree from the Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing, China, in 2022. He is currently an associate professor with the Aerospace Information Research Institute, Chinese Academy of Sciences. His research interests include spatial intelligence, embodied intelligence, and swarm intelligence, concentrating on multi-agent learning, multimodal fusion, 3D scene understanding and spatiotemporal data analysis.
[Uncaptioned image] Yuanchang Yue received the B.Sc. degree from jiangnan University, wuxi, China, in 2022. He is currently a master’s student with the Aerospace Information Research Institute, Chinese Academy of Sciences. His research interests include embodied intelligence, and task planning.
[Uncaptioned image] Youzhi Liu received the B.Sc. degree from Hunan University, changsha, China, in 2022. He is currently a Ph.D student with the Aerospace Information Research Institute, Chinese Academy of Sciences. His research interests include embodied intelligence, and visual language navigation.
[Uncaptioned image] Xian Sun received the B.Sc. degree from the Beijing University of Aeronautics and Astronautics, Beijing, China, in 2004, and the M.Sc. and Ph.D. degrees from the Institute of Electronics, Chinese Academy of Sciences, Beijing, in 2009. He is a Professor with the Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing, China. His research interests include computer vision, geospatial data mining, and remote sensing image understanding.
[Uncaptioned image] Kun Fu received the B.Sc., M.Sc., and Ph.D. degrees from the National University of Defense Technology, Changsha, China, in 1995, 1999, and 2002, respectively. He is a Professor with the Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing, China. His research interests include computer vision, remote sensing image understanding, geospatial data mining, and visualization.